OpenStudy (anonymous):

what is the sum of the finite arithmetic series? (-5)+0+5+10+....+65 900.A) 455.B) 450.C) 445.D)

OpenStudy (anonymous):

Okay, well, the sum of an arithmetic series is \[\frac{n(a_1+a_n)}{2}=S_n\] Where n is the number of terms in the series, a(1) is the first term and a(n) is the nth term.

OpenStudy (anonymous):

nth term?

OpenStudy (anonymous):

i.e. the number of terms you want to find. So if you only wanted -5, 0, 5, 10, 15 n=5 and a(n) would be 15 (the 5th term in the sequence)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Okay, so you know a(1) and a(n) right? Now you just need to determine how many terms are in that sequence from -5 until 65. So, to do that you'll want to know how much you are adding to the series each time.

OpenStudy (anonymous):

which is 5

OpenStudy (anonymous):

Yes. So, how can you use that to find the number of terms in the sequence from -5 to 65?

OpenStudy (anonymous):

add it, from 15 its 11 to get to 65 but then what

OpenStudy (anonymous):

Plug the number of terms you get into the equation for n. -5 for a(1) and 65 for a(n) And a hint for arithmetic sequences since it is always adding the same number each time, you can just divide the n-th term by the number you add each time if you have 0 in the set. So 65/5 = 13 (that would be all the terms from 5 to 65) then just add 2 more to that for 0 and -5